Write the equation of the line that contains the indicated pair of points. Express the final equation in standard form.
(2, 1), (11, 10)
1
Expert's answer
2020-02-24T09:40:07-0500

In the given question there are two points indicated, take them as A(2,1) and B (11,10),

so we have to find the equation of line passing through these points.

We know the general formula of equation of line passing through two point


(y-y1)= "\frac{ y_2-y_1}{x_2-x_1} (x-x_1)"


here "x_1= 2, x_2= 11, y_1=1,y_2=10"

now put these values in the above equation,

we get

"(y-1) = \frac{10-1}{11-2} (x-2)"

(y-1)= "\frac{9}{9} (x-2)"

(y-1)=1(x-2)

y-1=x-2 is equation of the line, but we have to represent it in the general form

and we know that general equation of line is y= mx+c.

Here, m= slope and c= intercept,

so, we can arrange it in general form

y-1= x-2

y= x-2+1

y=x-1 is standard equation of the line.